A particle performs linear SHM of amplitude 5 cm at what displacement from its mean position will the potential energy of the particle be 20% of the total energy
1
Expert's answer
2021-01-13T11:36:50-0500

The total energy is


"T=\frac12kA^2."

The potential energy corresponding to a position that results in being 20% of the total energy is


"P=\frac12kd^2=0.2T.\\\space\\
\frac12kd^2=0.2\cdot\frac12kA^2,\\\space\\
d^2=0.2A^2,\\
d=A\sqrt{0.2}=5\sqrt{0.2}=2.24\text{ cm}."